# Riemannian Flow Matching for Brain Connectivity Matrices via Pullback Geometry

### v1

@@ Appendix C Proof of Proposition 1: Riemannian CFM loss function on pullback manifolds
Furthermore, since $(\phi\circ\phi^{-1})(z)=z$, we have $\operatorname{D}\phi(\phi^{-1}(z))\circ\operatorname{D}\phi^{-1}(z)=\text{Id}_{E}$. Hence, we get
@@ Appendix E Proof of Proposition 3: Equivalence of Runge–Kutta iterates
$\dot{x}(t)$ $=u_{\theta}^{\mathcal{M}}(t,x(t),y)\quad\text{on manifold }\mathcal{M}$
@@ Appendix F The affine-invariant metric for correlation matrices
Appendix G Comparison of Baseline Methods: Assumptions, Strengths, and Limitations
  • TriangDDPM or TriangCFM: They are deep generative models that achieve strong performance across both CAS and quality-based metrics. However, their outputs are limited to triangular entities, requiring a post-processing projection step to obtain valid positive definite matrices. Unfortunately, this step alters the eigenvalue spectrum significantly, which degrades the quality and structure of the generated matrices (see Table 2 and Appendix L).
  • Riemannian CFM: This is a foundational contribution in the family of flow matching methods, providing a general theoretical framework for generative modeling on Riemannian manifolds. However, its current implementations are limited to a few specific manifolds such as spheres and tori, which are not directly applicable to our task. As a result, we adopt an alternative approach (i.e., pullback geometry) to enable generative modeling on SPD and correlation manifolds.
  • DiffeoCFM: The proposed approach introduces a novel Riemannian CFM framework using the pullback geometry, enabling direct data generation on SPD and correlation manifolds, particularly helpful in our neuroscience and neuroengineering tasks. It trains efficiently (see Figure 3), achieves strong performance on both quality and CAS metrics (see Table 1), and consistently demonstrates neurophysiological plausibility across two distinct evaluation protocols in both fMRI and EEG (see Section 5.2).
Appendix H Training time comparison
Figure 3: Training efficiency on ${\mathbb{S}_{d}^{++}}$ (simulated data). Log-scale training time over $1000$ epochs for generating SPD matrices. We vary either the number of samples $n$ or the matrix dimension $d$, keeping the other fixed. DiffeoCFM leverages the diffeomorphism $\phi_{\mathbb{S}_{d}^{++}}$ (9) and is at least 1000×1000$\times$1000 × faster than SPD-DDPM [38]. Importantly, DiffeoCFM can also generate natively correlation matrices (${\text{Corr}_{d}}$), and it performs with same training time as on ${\mathbb{S}_{d}^{++}}$.
Appendix I Datasets and preprocessing
Resting-state fMRI datasets and preprocessing
@@ Appendix I Datasets and preprocessing
Sessions with missing or non-standard protocol (e.g., 2C in BNCI 2015-001) were discarded. We further filtered time series whose covariance matrices have high values (max entry $>10^{4}$) or statistical outliers based on Mahalanobis distance from the group mean. This retained the top 90% most consistent trials for downstream modeling.
@@
BNCI2014-002: The BNCI2014-002111https://neurotechx.github.io/moabb/generated/moabb.datasets.BNCI2014_002.html dataset, provided through the BNCI Horizon 2020 initiative, includes recordings from 13 subjects engaged in motor imagery tasks. Participants were instructed to imagine movements of either their right hand or both feet for a sustained duration of 5 seconds, guided by visual cues. EEG signals were captured using an amplifier and active Ag/AgCl electrodes at a sampling rate of 512 Hz, with a total of 15 electrode channel applied to each subject. The experimental protocol comprised eight sessions per participant, each including 80 trials for hand and foot imagery, summing to 160 trials. EEG epochs were extracted from 3.0 to 8.0 seconds relative to cue onset, aligning with the motor imagery window.
@@
BNCI2015-001: The BNCI2015-001222https://neurotechx.github.io/moabb/generated/moabb.datasets.BNCI2015_001.html dataset, released as part of the BNCI Horizon 2020 project, contains EEG recordings from 12 individuals who performed motor imagery tasks involving either the right hand or both feet. EEG data were acquired at a 512 Hz sampling rate, preprocessed with a 0.5–100 Hz bandpass filter and a 50 Hz notch filter. Each trial spanned 5 seconds, beginning at 3.0 seconds after cue onset and ending at 8.0 seconds, corresponding to the motor imagery phase. The dataset provides EEG data from 13 electrode channels arranged in the following order: FC3, FCz, FC4, C5, C3, C1, Cz, C2, C4, C6, CP3, CPz, and CP4. For most participants (Subjects 1–8), recordings were conducted over two sessions on consecutive days. In contrast, Subjects 9–12 completed three sessions. Each session consisted of 100 trials per class, resulting in 200 trials per session per subject.
Appendix J Implementations
Evaluation metrics
We report two families of metrics: quality and Classification Accuracy Score (CAS) metrics. To assess quality, we compute $\alpha$-precision and $\beta$-recall [1] using the EvaGeM library. To assess classification utility, we follow the CAS protocol [53]: we train a classifier on generated samples and evaluate it on real data, reporting ROC-AUC and F1 scores. Classification is performed using logistic regression after tangent space mapping with the affine-invariant metric. We use logistic regression with 5-fold cross-validation to select the inverse regularization strength $C$ from the grid $\{10^{-4},10^{-3},\dotsc,10^{4}\}$. The logistic regression model uses the liblinear solver and balances class weights.
Deep learning model and training/sampling setups
TriangDDPM, TriangCFM and DiffeoCFM employ a two-layer MLP with $512$ hidden units, trained using AdamW [42] with a learning rate of $10^{-3}$ and batch size $64$. Training runs for $200$ epochs on fMRI and $2000$ epochs on EEG. Time integration follows the midpoint method with a step size of $10^{-3}$.
Appendix K Constraints Satisfaction (${\mathbb{S}_{d}^{++}}$ and ${\text{Corr}_{d}}$)
@@
Appendix M DiffeoCFM on EEGs
M.1 Filter’s Topographic Map
@@
Appendix N DiffeoCFM on fMRI
N.1 fMRI Connectome Plotting
@@
We use Nilearn [^3] https://nilearn.github.io/stable/index.html for neuroimaging analysis and visualization based on the MSDL brain atlas. The atlas is retrieved using the fetch__\__atlas__\__msdl function, which provides probabilistic maps of functional brain networks in the form of a 4D NIfTI image. The maps attribute extracts this image, while find__\__probabilistic__\__atlas__\__cut__\__coords computes representative 3D coordinates for each network, typically corresponding to the spatial center of the activation, to support anatomical localization. In addition, predefined region labels are extracted to annotate the corresponding brain areas. These components together enable the visualization of brain networks, including plotting the regions with spatial coordinates and labels.
The connectome plots are grouped into two categories: "CN" and "non-CN". The adjacency__\__matrix used in nilearn.plotting.plot__\__connectome is computed as the Fréchet mean of the corresponding group of connectivity matrices in all the 5-fold experiments, and all edge thresholds in the plots are set to 90%.
@@
Appendix O Plotting of Generated Samples in Real Data Neighborhoods
Figure 9-14 present the selected generated fMRI samples that are nearest to real data points in terms of Frobenius distance, categorized by "CN" and "non-CN" from the ADNI, ABIDE, and OASIS3 datasets.. Figure 15-18 show the selected generated EEG samples that are nearest to real samples in Frobenius distance, separately for hand and feet motor imagery tasks from the BNCI2014-002 and BNCI2015-001 datasets. Each real sample is annotated with its index in the original dataset. The first, second, and third columns correspond to samples generated by DiffeoGauss, TriangCFM, and DiffeoCFM, respectively. The numeric values on the generated samples indicate the Frobenius distance to the corresponding real sample. Each generated sample shown is the nearest one (in Frobenius distance) to the real sample within its class. Due to its consistently highest recall across multiple datasets, DiffeoCFM is more likely to produce generated samples that are closer to real samples, i.e., with the smallest Frobenius distance, compared to those generated by other methods. The matrices generated by TriangCFM shown in the figures are presented without applying the projection step. Before applying the projection, the outputs of TriangCFM already exhibit strong recall scores.

### v2

@@ Appendix C Proof of Proposition 1: Riemannian CFM loss function on pullback manifolds
Furthermore, since $(\phi\circ\phi^{-1})(z)=z$, we have $\operatorname{D}\phi(\phi^{-1}(z))\circ\operatorname{D}\phi^{-1}(z)=\text{Id}_{E}$. qHence, we get
@@ Appendix E Proof of Proposition 3: Equivalence of Runge–Kutta iterates
[…]
@@ Appendix F The affine-invariant metric for correlation matrices
Appendix G Datasets and Preprocessing
Resting‐state fMRI datasets and preprocessing
@@ Appendix G Datasets and Preprocessing
Sessions with missing or non-standard protocol (e.g., 2C in BNCI2015-001) were discarded. We further filtered time series whose covariance matrices have high values (max entry $10^{4}$) or statistical outliers based on Mahalanobis distance from the group mean. This retained the top 90% most consistent trials for downstream modeling.
@@
BNCI2014-002: The BNCI2014-002333https://neurotechx.github.io/moabb/generated/moabb.datasets.BNCI2014_002.html dataset, provided through the BNCI Horizon 2020 initiative, includes recordings from 13 subjects engaged in motor imagery tasks. Participants were instructed to imagine movements of either their right hand or both feet for a sustained duration of 5 seconds, guided by visual cues. EEG signals were captured using an amplifier and active Ag/AgCl electrodes at a sampling rate of 512 Hz, with a total of 15 electrode channel applied to each subject. The experimental protocol comprised eight sessions per participant, each including 80 trials for hand and foot imagery, summing to 160 trials. EEG epochs were extracted from 3.0 to 8.0 seconds relative to cue onset, aligning with the motor imagery window.
@@
BNCI2015-001: The BNCI2015-001444https://neurotechx.github.io/moabb/generated/moabb.datasets.BNCI2015_001.html dataset, released as part of the BNCI Horizon 2020 project, contains EEG recordings from 12 individuals who performed motor imagery tasks involving either the right hand or both feet. EEG data were acquired at a 512 Hz sampling rate, preprocessed with a 0.5–100 Hz bandpass filter and a 50 Hz notch filter. Each trial spanned 5 seconds, beginning at 3.0 seconds after cue onset and ending at 8.0 seconds, corresponding to the motor imagery phase. The dataset provides EEG data from 13 electrode channels arranged in the following order: FC3, FCz, FC4, C5, C3, C1, Cz, C2, C4, C6, CP3, CPz, and CP4. For most participants (Subjects 1–8), recordings were conducted over two sessions on consecutive days. In contrast, Subjects 9–12 completed three sessions. Each session consisted of 100 trials per class, resulting in 200 trials per session per subject.
Appendix H Comparison of Baseline Methods: Assumptions, Strengths, and Limitations
  • TriangDDPM or TriangCFM: They are deep generative models that achieve strong performance across both CAS and quality-based metrics. However, their outputs are limited to triangular entities, requiring a post-processing projection step to obtain valid positive definite matrices. Unfortunately, this step alters the eigenvalue spectrum significantly, which degrades the quality and structure of the generated matrices (see Table 2 and Appendix M).
  • Riemannian CFM: This is a foundational contribution in the family of flow matching methods, providing a general theoretical framework for generative modeling on Riemannian manifolds. However, its construction is geometry-specific and requires a dedicated implementation for each manifold. Furthermore, on SPD matrices, we observe that it is roughly $8\times$ slower to train and $10\times$ slower to sample from than DiffeoCFM (see Table 1), due to the repeated computation of affine-invariant geodesics, their derivatives and Riemannian norms.
  • DiffeoCFM: The proposed approach introduces a novel Riemannian CFM framework using the pullback geometry, enabling direct data generation on SPD and correlation manifolds, particularly helpful in neuroscience and neuroengineering tasks. It trains and sampels efficiently (see Figure 4 and Table 1), achieves strong performance on both quality and CAS metrics (see Table 1), and consistently demonstrates neurophysiological plausibility across two distinct evaluation protocols in both fMRI and EEG (see Section 5.2).
Appendix I Summary Figure
While Table 1 provides a comprehensive numerical evaluation of all methods, a visual representation can offer clearer insights into the practical trade-offs between model performance and computational efficiency. To this end, Figure 3 plots a unified performance metric,the Average F1 Score, against both training and sampling times. This allows for a direct comparison of the performance-cost profile of each method for both fMRI and EEG data.
Figure 3: Trade-off between generative performance and computational cost for fMRI (top) and EEG (bottom) data. The figure plots the Average F1 Score against Training Time (left) and Sampling Time (right). The Average F1 Score is the mean of the quality metric ($\alpha,\beta$-F1) and the CAS F1-score from Table 1. Each point marks the mean performance across all splits and datasets for a given modality, with error bars and shaded regions indicating the standard deviation. The dashed gray line represents the Real Data baseline, which serves as an empirical upper bound. Time is shown on a logarithmic scale.
Appendix J Scalability and Computational cost of SPD–DDPM [45] vs DiffeoCFM
Figure 4: Training efficiency on $\mathbb{S}_{d}^{++}$} (simulated data). Log-scale training time over $1000$ epochs for generating SPD matrices. We vary either the number of samples $n$ or the matrix dimension $d$, keeping the other fixed. DiffeoCFM leverages the diffeomorphism $\phi_{\mathbb{S}_{d}^{++}}$ (9) and is at least $1000\times$ faster than SPD-DDPM [45]. Importantly, DiffeoCFM can also generate natively correlation matrices ($\text{Corr}_{d}$}), and it performs with same training time as on $\mathbb{S}_{d}^{++}$}.
Appendix K Ablation Study
We conduct an ablation study to analyze the sensitivity of our proposed DiffeoCFM model to its neural network architecture. The experiments are performed on the ABIDE dataset using the Cholesky diffeomorphism. We investigate two key architectural hyperparameters: the width (size of hidden layers) and the depth (number of hidden layers).
First, we fix the network depth to a single hidden layer and vary its width from 8 to 1024, as detailed in Table 3. Second, we fix the layer width to 512 neurons and vary the depth from one to four hidden layers, with results presented in Table 4. For both experiments, performance is evaluated using the $\alpha,\beta$-F1 score, which is the harmonic mean of $\alpha$-precision and $\beta$-recall. We also report the average training and sampling times over 10 splits.
The results from Table 3 indicate that performance generally improves with network width, peaking at 256 and 512 hidden units. A wider network of 1024 units shows a drop in performance. Regarding network depth, Table 4 clearly shows that a single hidden layer achieves the best results. Deeper models exhibit a consistent decline in performance. Based on these findings, a single hidden layer with a width of 512 was chosen for our main experiments, as it offers the best performance without unnecessary complexity.
Table 3: Ablation study on the width of the hidden layer for the DiffeoCFM model on the ABIDE dataset. The network depth is fixed to one hidden layer. We report the mean and standard deviation of the $\alpha,\beta$-F1 score, training time, and sampling time across 10 splits. Best values are in bold.
Hidden dim. $\alpha,\beta$-F1 Training time (s.) Sampling time (s.)
8 0.47 ± 0.06 29.07 ± 1.07 0.35 ± 0.01
16 0.47 ± 0.09 27.46 ± 1.24 0.29 ± 0.03
32 0.54 ± 0.05 28.36 ± 1.26 0.26 ± 0.02
64 0.57 ± 0.07 29.69 ± 1.31 0.30 ± 0.03
128 0.60 ± 0.05 31.01 ± 1.32 0.35 ± 0.04
256 0.63 ± 0.07 31.39 ± 0.96 0.32 ± 0.04
512 0.63 ± 0.06 31.36 ± 1.14 0.31 ± 0.02
1024 0.49 ± 0.12 32.34 ± 1.17 0.36 ± 0.05
Table 4: Ablation study on the depth of the neural network for the DiffeoCFM model on the ABIDE dataset. The width of all hidden layers is fixed to 512. We report the mean and standard deviation of the $\alpha,\beta$-F1 score, training time, and sampling time across 10 splits. Best values are in bold.
Depth $\alpha,\beta$-F1 Training time (s.) Sampling time (s.)
1 0.62 ± 0.05 33.06 ± 1.05 0.38 ± 0.03
2 0.52 ± 0.09 32.93 ± 1.00 0.36 ± 0.02
3 0.46 ± 0.05 34.36 ± 1.00 0.46 ± 0.04
4 0.50 ± 0.08 35.68 ± 1.01 0.48 ± 0.08
Appendix L Constraint Satisfaction ($\mathbb{S}_{d}^{++}$} and $\text{Corr}_{d}$})
@@ Appendix I Summary Figure
Appendix N DiffeoCFM on EEGs
N.1 Filter’s Topographic Map
@@
Appendix O DiffeoCFM on fMRI
O.1 fMRI Connectome Plotting
@@
We use Nilearn [14] for neuroimaging analysis and visualization based on the MSDL brain atlas. The atlas is retrieved using the fetch_\_atlas_\_msdl function, which provides probabilistic maps of functional brain networks in the form of a 4D NIfTI image. The maps attribute extracts this image, while find_\_probabilistic_\_atlas_\_cut_\_coords computes representative 3D coordinates for each network, typically corresponding to the spatial center of the activation, to support anatomical localization. In addition, predefined region labels are extracted to annotate the corresponding brain areas. These components together enable the visualization of brain networks, including plotting the regions with spatial coordinates and labels.
The connectome plots are grouped into two categories: "CN" and "non-CN". The adjacency_\_matrix used in nilearn.plotting.plot_\_connectome is computed as the Fréchet mean of the corresponding group of connectivity matrices in all the 5-fold experiments, and all edge thresholds in the plots are set to 90%.
@@
Appendix P Plotting of Generated Samples in Real Data Neighborhoods
Figure 10-15 present the selected generated fMRI samples that are nearest to real data points in terms of Frobenius distance, categorized by "CN" and "non-CN" from the ADNI, ABIDE, and OASIS3 datasets.. Figure 16-19 show the selected generated EEG samples that are nearest to real samples in Frobenius distance, separately for hand and feet motor imagery tasks from the BNCI2014-002 and BNCI2015-001 datasets. Each real sample is annotated with its index in the original dataset. The first, second, and third columns correspond to samples generated by DiffeoGauss, TriangCFM, and DiffeoCFM, respectively. The numeric values on the generated samples indicate the Frobenius distance to the corresponding real sample. Each generated sample shown is the nearest one (in Frobenius distance) to the real sample within its class. Due to its consistently highest recall across multiple datasets, DiffeoCFM is more likely to produce generated samples that are closer to real samples, i.e., with the smallest Frobenius distance, compared to those generated by other methods. The matrices generated by TriangCFM shown in the figures are presented without applying the projection step. Before applying the projection, the outputs of TriangCFM already exhibit strong recall scores.